Consider the problem of solving xlog5(x)=136. Formulate this as a problem of solving a nonlinear equation f(x)=0.

      f(x)=       
Answer #1

alag, (3)=136. lag, (*)= 136 136 n 5 X = Let fere = 5 136 x -2 u. 136 This fex) = 5 - О. 136 a 136 2 fexz0 35 -n=0 fx=5 na 13